# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### v1

@@ F.1 Canonical Twistors and Dimensional Reduction
$\mu(z)=\mathrm{i}e^{-\mathrm{i}\pi/4}z^{1/2}\begin{pmatrix}1\\ -z\end{pmatrix}$ (F.1)
@@
The defining geometric characteristic of this canonical configuration is its behavior under the symmetric Pauli matrices. The spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$. Evaluating the symmetric bilinear for the first spatial component:
@@
$=\mathrm{i}\left[\left(\frac{1}{z}\right)(-z)+(1)(1)\right]=\mathrm{i}(-1+1)=0$ (F.2)
Because the tangent vector component $f^{\prime}_{1}$ vanishes identically everywhere on the worldsheet, the target space coordinate $X_{1}$ is strictly constant. This exact algebraic constraint dynamically crushes the generic 4-dimensional minimal surface into a 3-dimensional hyperplane, forming the exact topological skeleton of the macroscopic 3D Helicoid.
F.2 The Euclidean Metric and Double Wick Rotation
@@
$\exp\left(2\rho\right)=R^{2}\frac{(1+|z|^{2})^{2}}{|z|^{2}}=R^{2}\left(|z|+\frac{1}{|z|}\right)^{2}$ (F.4)
@@
$\exp\left(2\rho\right)=R^{2}\left(e^{\sigma_{E}}+e^{-\sigma_{E}}\right)^{2}=4R^{2}\cosh^{2}(\sigma_{E})$ (F.5)
@@
$ds^{2}_{E}=\exp\left(2\rho\right)\frac{dzd\bar{z}}{|z|^{2}}=4R^{2}\cosh^{2}(\sigma_{E})\left(d\sigma_{E}^{2}+d\tau_{E}^{2}\right)$ (F.6)
@@
$\sigma_{E}=\mathrm{i}\sigma_{M}\implies d\sigma_{E}^{2}=-d\sigma_{M}^{2}$ (F.7)
Under this transformation, the hyperbolic cosine strictly rotates into the trigonometric cosine: $\cosh(\mathrm{i}\sigma_{M})=\cos(\sigma_{M})$. Extracting an overall minus sign to reflect the intrinsic Lorentzian (,+)(-,+) worldsheet signature, the exact Minkowski metric is:
$ds^{2}_{M}=4R^{2}\cos^{2}(\sigma_{M})\left(d\tau_{E}^{2}-d\sigma_{M}^{2}\right)$ (F.8)
@@
$dr=\cos(\sigma_{M})d\sigma_{M}\implies dr^{2}=\cos^{2}(\sigma_{M})d\sigma_{M}^{2}$ (F.9)
@@
$ds^{2}_{M}=4R^{2}\left[(1-r^{2})d\tau_{E}^{2}-dr^{2}\right]$ (F.10)
This flawlessly reproduces the canonical Nambu-Goto metric $\propto(1-r^{2})dt^{2}-dr^{2}$ for a macroscopic string rotating at the speed of light.
F.3 Action Evaluation and the Lüscher Shift
Finally, we pedantically evaluate the total effective action over the fundamental domain of this Minkowski metric. The determinant of the conformal Minkowski metric $ds^{2}_{M}=4R^{2}\cos^{2}(\sigma_{M})(d\tau_{E}^{2}-d\sigma_{M}^{2})$ is $g=16R^{4}\cos^{4}(\sigma_{M})$, yielding the exact proper area measure:
$\sqrt{-g}=4R^{2}\cos^{2}(\sigma_{M})$ (F.11)
@@ F.3 Action Evaluation and the Lüscher Shift
Area $=\int_{0}^{2\pi}d\tau_{E}\int_{0}^{\pi/2}4R^{2}\cos^{2}(\sigma_{M})d\sigma_{M}$
$=(2\pi)(4R^{2})\left(\frac{\pi}{4}\right)=2\pi^{2}R^{2}$ (F.12)
Simultaneously, the Liouville action dictates that the worldsheet carries the kinetic anomaly $\frac{1}{12\pi}|\nabla\rho|^{2}$. Because the kinetic anomaly density shares the exact same spatial singularity as the conformal area density at the twistor pole, it mathematically merges into the classical area term. As established in the main text, it acts strictly as a rigid, local shift to the effective string tension:
$\sigma_{\text{eff}}=\sigma+\frac{1}{12\pi R^{2}}$ (F.13)
Multiplying the classical Minkowski area by this effective tension, we obtain the complete 1-cycle bulk action:
@@
$=-\left(\sigma+\frac{1}{12\pi R^{2}}\right)\left(2\pi^{2}R^{2}\right)$
$=-2\pi^{2}\sigma R^{2}-\frac{\pi}{6}$ (F.14)
This rigorous derivation explicitly demonstrates that the universally mandated $\pi/6$ Lüscher phase is identically generated by the exact geometric mapping of the twistor string to the Minkowski cylinder.

### v2

@@ D.1 Chiral Twistors and Dimensional Breakout
$\mu(z)=\mathrm{i}e^{-\mathrm{i}\pi/4}z^{1/2}\begin{pmatrix}1\\ 0\end{pmatrix}$ (D.1)
@@
The defining geometric characteristic of this configuration is its behavior under the symmetric Pauli matrices. The spatial tangent vectors of the minimal surface are constructed via $f^{\prime}_{k}\propto\lambda^{T}\sigma_{k}\mu$. Evaluating the symmetric bilinear for the first spatial component:
@@
$=\left(e^{\mathrm{i}\pi/4}z^{-3/2}\right)(0)+\left(e^{\mathrm{i}\pi/4}z^{-1/2}\right)\left(\mathrm{i}e^{-\mathrm{i}\pi/4}z^{1/2}\right)=\mathrm{i}$ (D.2)
Unlike the generic configuration where this component vanished identically everywhere on the worldsheet, here the tangent vector component $f^{\prime}_{1}$ evaluates to a non-zero constant. Consequently, the target space coordinate $X_{1}$ is no longer strictly constant. This permanently breaks the exact algebraic constraint that previously crushed the generic minimal surface into a 3-dimensional hyperplane, meaning this chiral surface genuinely spans a 4-dimensional embedding.
D.2 The Euclidean Metric and Parity Projection
@@
$\exp\left(2\rho\right)=R^{2}\frac{1+|z|^{2}}{|z|^{2}}=R^{2}\left(1+\frac{1}{|z|^{2}}\right)$ (D.4)
@@
$\exp\left(2\rho\right)=R^{2}\left(1+e^{-2\sigma_{E}}\right)$ (D.5)
@@
$ds^{2}_{E}=\exp\left(2\rho\right)(d\sigma_{E}^{2}+d\tau_{E}^{2})$
$=R^{2}\left(1+e^{-2\sigma_{E}}\right)\left(d\sigma_{E}^{2}+d\tau_{E}^{2}\right)$ (D.6)
@@
$\sigma_{E}=\mathrm{i}\sigma_{M}\implies d\sigma_{E}^{2}=-d\sigma_{M}^{2}$ (D.7)
Under this transformation, the exponential strictly rotates into a complex phase: $e^{-2\mathrm{i}\sigma_{M}}=\cos(2\sigma_{M})-\mathrm{i}\sin(2\sigma_{M})$.
Because the macroscopic QCD string is strictly parity-conserving, the physical state is evaluated as the parity-even arithmetic mean of the left-chiral and right-chiral target space configurations. The conjugate right-chiral surface identically provides the inverse complex phase $e^{+2\mathrm{i}\sigma_{M}}$. Taking their arithmetic mean rigorously annihilates the imaginary sine terms, strictly isolating the real, physical part of the metric:
$\text{Re}\left[\exp\left(2\rho\right)\right]=R^{2}\big(1+\cos(2\sigma_{M})\big)=2R^{2}\cos^{2}(\sigma_{M})$ (D.8)
Extracting an overall minus sign to reflect the intrinsic Lorentzian (,+)(-,+) worldsheet signature, the exact physical Minkowski metric is:
$ds^{2}_{M}=2R^{2}\cos^{2}(\sigma_{M})\left(d\tau_{E}^{2}-d\sigma_{M}^{2}\right)$ (D.9)
@@
$dr=\cos(\sigma_{M})d\sigma_{M}\implies dr^{2}=\cos^{2}(\sigma_{M})d\sigma_{M}^{2}$ (D.10)
@@
$ds^{2}_{M}=2R^{2}\left[(1-r^{2})d\tau_{E}^{2}-dr^{2}\right]$ (D.11)
This flawlessly reproduces the canonical Nambu-Goto metric $\propto(1-r^{2})d\tau_{E}^{2}-dr^{2}$ for a macroscopic string rotating at the speed of light, but exactly halved in overall magnitude relative to the generic three-dimensional configuration.
D.3 Action Evaluation and the Halved Lüscher Shift
Finally, we pedantically evaluate the total effective action over the fundamental domain of this purely real Minkowski metric. The determinant of the conformal Minkowski metric $ds^{2}_{M}=2R^{2}\cos^{2}(\sigma_{M})(d\tau_{E}^{2}-d\sigma_{M}^{2})$ yields the exact proper area measure:
$\sqrt{-g}=2R^{2}\cos^{2}(\sigma_{M})$ (D.12)
@@ D.3 Action Evaluation and the Halved Lüscher Shift
Area $=\int_{0}^{2\pi}d\tau_{E}\int_{0}^{\pi/2}2R^{2}\cos^{2}(\sigma_{M})d\sigma_{M}$
$=(2\pi)(2R^{2})\left(\frac{\pi}{4}\right)=\pi^{2}R^{2}$ (D.13)
Simultaneously, the Liouville action dictates that the worldsheet carries the kinetic anomaly $\frac{1}{12\pi}|\nabla\rho|^{2}$. Because the parity-symmetric conformal factor shares the exact same spatial singularity at the twistor pole as the generic string, it mathematically merges into the classical area term with the identical anomaly coefficient. As established in the main text, it acts strictly as a rigid, local shift to the effective string tension:
$\sigma_{\text{eff}}=\sigma+\frac{1}{12\pi R^{2}}$ (D.14)
Multiplying the purely real classical Minkowski area by this effective tension, we obtain the complete one-cycle bulk action:
@@
$=-\left(\sigma+\frac{1}{12\pi R^{2}}\right)\left(\pi^{2}R^{2}\right)$
$=-\pi^{2}\sigma R^{2}-\frac{\pi}{12}$ (D.15)
This rigorous derivation explicitly demonstrates that for the parity-projected chiral string, the classical string area is exactly halved, which identically matches the universally mandated topological Lüscher phase $\pi/12$, though from a completely different physical and mathematical origins, which we discuss in the text.
Appendix E The Pinch Singularity and the Topological Barrier
To mathematically illustrate the appearance of the conjugate pole at $1/\bar{z}_{0}$ and the resulting impenetrable topological barrier at the unit circle, consider the evaluation of the Liouville kinetic anomaly term in the effective action. In for a meromorphic twistor field with one pole $z_{0}$, the conformal factor generates an area integral of the form:
$I=\int_{D}d^{2}z\frac{1}{(z-z_{0})(\bar{z}-\bar{z}_{0})}$ (E.1)
We can rewrite the integrand as a total derivative with respect to $\bar{z}$:
$I\propto\int_{D}dzd\bar{z}\,\partial_{\bar{z}}\left(\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}\right)$ (E.2)
By applying Stokes’ theorem, the two-dimensional bulk integral over the disk $D$ reduces exactly to a one-dimensional contour integral over the boundary unit circle $|z|=1$:
$I\propto\oint_{|z|=1}dz\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}$ (E.3)
Crucially, on the unit circle, the complex conjugate is strictly given by the geometric inversion $\bar{z}=1/z$. Substituting this into the logarithm yields the exact boundary representation of the action:
$I\propto\oint_{|z|=1}dz\frac{\log(1/z-\bar{z}_{0})}{z-z_{0}}$ (E.4)
This exact formula explicitly reveals the analytic structure governing the boundary dynamics. The integrand possesses a simple pole at $z=z_{0}$ (originating from the twistor itself) and a logarithmic branch point at $z=1/\bar{z}_{0}$ (originating from the boundary inversion of the conjugate twistor).
Because the physical domain restricts the pole to the interior of the disk ($|z_{0}|<1$), its conjugate mirror strictly resides outside the disk ($|1/\bar{z}_{0}|>1$). However, if any continuous Langevin fluctuation attempts to push the twistor pole across the boundary ($|z_{0}|\to 1$), its conjugate mirror must simultaneously approach the boundary from the outside.
At the exact moment of crossing, these two singularities violently collide directly on the physical integration contour $|z|=1$, trapping the integration path between them. This creates a severe, non-integrable pinch singularity, causing the action integral to macroscopically diverge. This geometric divergence acts as an infinite energetic barrier, rigorously protecting the twistor pole from escaping and locking the string into its quantized topological sector.
Appendix F Algebraic Derivation of Spin Holonomy and Trace T
To rigorously justify the topological selection rules governing the meson spectrum, we compute the path-ordered Dirac trace $T$ for the one-pole sector ($z_{0}=0$). The fundamental link matrix $\mathcal{M}(\theta)$ at the boundary $|z|=1$ is defined as:
$\mathcal{M}(\theta)=i\mathcal{P}(\theta)=-2\left(e^{i\theta}\eta\xi^{\dagger}-e^{-i\theta}\xi\eta^{\dagger}\right)$ (F.1)
where $\eta$ is the basis spinor of the frozen $SU(2)_{R}$ sector and $\xi(\theta)$ is the rotating spinor of the active $SU(2)_{L}$ sector. In the basis where $\eta=\begin{pmatrix}1\\ 0\end{pmatrix}$ and the twistor $\lambda(z)\propto 1/z$ implies $\xi(\theta)=\begin{pmatrix}0\\ e^{-i\theta}\end{pmatrix}$ at the boundary, the local matrix evaluates to:
$\mathcal{M}(\theta)=\begin{pmatrix}0&-2e^{2i\theta}\\ 2e^{-2i\theta}&0\end{pmatrix}$ (F.2)
This matrix represents a local $SU(2)$ rotation that flips the chiral state. The propagator $\mathcal{U}$ for one physical cycle is the path-ordered product of these matrices. For two adjacent steps $\theta_{1}$ and $\theta_{2}=\theta_{1}+\Delta\theta$, we find:
$\mathcal{M}(\theta_{2})\mathcal{M}(\theta_{1})=\begin{pmatrix}-4e^{2i\Delta\theta}&0\\ 0&-4e^{-2i\Delta\theta}\end{pmatrix}$ (F.3)
Assuming even number $2L$ of points on a unit circle in discretized loop, and multiplying these matrices around the entire loop $\oint d\theta=2\pi$ reduces the propagator for one cycle to:
$\mathcal{U}=(-4)^{L}\begin{pmatrix}e^{4\pi i}&0\\ 0&e^{-4\pi i}\end{pmatrix}$ (F.4)
The physical trace $T$
[…]
The real factor renormalizes the bare mass of quark, so we drop it, but the phase is interesting, as ir accumulkates in the windings and leads to the negative sign in front of $\exp\left(\Delta S\right)$. which changes the quantization condition by selecting odd or even n in B-S quantization $\Delta S_{eff}=\mathrm{i}\pi n$ depending on the parity of $L$.